OpenStudy (amber15):

a topical cream is sold in a 26 gram container. the average amount of topical cream used per application is 2/7 grams . how many applications can be made with the container

OpenStudy (anonymous):

\[\frac{ 2 }{ 7 } \div \frac{ 26 }{ 1 }\] which becomes \[\frac{ 2 }{ 7 } \times \frac{ 1 }{ 26 } = ?\]

OpenStudy (amber15):

so that becomes 1/91

OpenStudy (anonymous):

That's correct